A gold club operator finds that she can sell 30 sets of golf clubs at $500 per set in a year. She predicts that for each $25 the decrease in price 3 extra sets of clubs could be sold. At what price should she sell the clubs to maximize the income?
Thanks so much in advance:)

Answer by lwsshak3(11628)

A gold club operator finds that she can sell 30 sets of golf clubs at $500 per set in a year. She predicts that for each $25 the decrease in price 3 extra sets of clubs could be sold. At what price should she sell the clubs to maximize the income?
***
let x=number of $25 decreases in price
revenue=price*quantity
R=(500-25x)(30+3x)
=15000-750x+1500x-75x^2
=15000-750x-75x^2
=-75x^2+750x+15000
divide by 75
=-x^2+10x+200
complete the square:
=-(x^2-10x+25)+25+200
R=-(x-5)^2+225
This is an equation of a parabola that opens down.
its vertex:(5,225)
number of $25 decreases in price=5
price should she sell the clubs to maximize the income=500-5*25=$375